OpenStudy (anonymous):

A 0.6 ml dose of a drug is injected into a patient steadily for one second. At the end of this time, the quantity, Q, of the drug in the body starts to decay exponentially at a continuous rate of 0.2% per second. Express Q as a continuous function of time t in seconds. Use the form given below. Q(t)= { A(t), 0 ≤ t<1 { B(t), 1

OpenStudy (anonymous):

so find A(t)

OpenStudy (anonymous):

how? I know t is 1 sec

OpenStudy (anonymous):

steadily injected means its a linear (line) equation at t=1 the dose=.6, at t=0 dose=0 (0,0) (1,.6)

OpenStudy (anonymous):

2 points so you can find a line

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

finding B(t) is a bit tricky and i really dont have the attention span atm to solve it we know every second it decreases by .2%

OpenStudy (anonymous):

but what equation am I plugging these numbers into?

OpenStudy (anonymous):

t=1 dose = .6 t=2 dose= .6(1-.002) t=3 dose=.6(1-.002)^2 can you make an equation for this?

OpenStudy (anonymous):

okay so after 2 second theres 0.998

OpenStudy (anonymous):

I still don't get what A(t) is

OpenStudy (anonymous):

you solve that yourself IMPORTANT LINE RELATED EQUATIONS TO KNOW AND MEMORIZE slope formula m= slope/ gradiant -- same thing \[m=\frac{y_2-y_1}{x_2-x_1}\] standard formula \[Ax+By=C\] point-slope formula \[y-y_1=m(x-x_1)\] slope-intercept formula b= y-intercept -- in the form of (0,y) \[y=mx+b\]

OpenStudy (anonymous):

Can someone tell me what A(t) is I can't figure it out??